40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1670     Entry time: Thu Jul 14 20:57:07 2016
Author: Antonio 
Type: DailyProgress 
Category: optic 
Subject: South path construction progress 

Summary
Following ID 1649 we start the implementation of the South cavity path.
Some optics has been placed along the path.

Conclusions 
A quick beam profile measurement and the implementation of the BB EOM
is the next step.
 


Things done:

  • Optics as shown in the setup have been placed;
  • References of the beam along the path have been aligned  (Those should be kept in place)


Notes:

  • The FI is mounted on a base which brings it NOT at 3" height; Two steering mirrors have been used to overcome this issue;
  • There is a second beam coming from the FI; I did not have a chance to play with it, but Andrew did and the origin of that is not totally understood;

Setup

Attachment 1: North_South_layout_progress_dist_final_copy.pdf  83 kB  | Hide | Hide all
North_South_layout_progress_dist_final_copy.pdf
ELOG V3.1.3-